# Tailored Design of Audio-Visual Speech Recognition Models using Branchformers

### v1

@@ III-A Model Architecture
$\boldsymbol{X}=PoswiseFFN\left(w_{audio}\boldsymbol{X}_{audio}+w_{video}\boldsymbol{X}_{video}\right)$ (4)
CTC/Attention Decoder. This state-of-the-art hybrid architecture [27] combines both the Markov assumptions of CTC [44] (properties in harmony with the speech nature) and the flexibility of the non-sequential alignments offered by the attention-based decoder [24]. As Figure 3 depicts, this module is composed of a linear projection as the CTC-based decoding branch and a 6-layer Transformer decoder of 4 attention heads of 64 dimensions attached to the Cross Entropy (CE) criterion. As proposed by [27], its loss function is defined as follows:
$\mathcal{L}=\alpha\log p_{ctc}(\boldsymbol{Y}|\boldsymbol{X})+(1-\alpha)\log p_{attn}(\boldsymbol{Y}|\boldsymbol{X})$ (5)
Language Model. In this work, we use the character-level Language Models (LM) publicly released by Ma et al. [6] for Spanish and English. These models, defined as 16-layer Transformer encoders [24], were estimated with text from different sources comprising more than 150M characters each.
$S=\lambda S_{ctc}+(1-\lambda)S_{attn}+\beta S_{lm}$ (6)
where $S_{ctc}$ and $S_{attn}$ are the scores of the CTC- and the Attention-based decoder, respectively, $\lambda$ is their corresponding relative weight, and $\beta$ and $S_{lm}$ refer to the LM decoding influence weight and the LM score, respectively.
III-B The Tailored Audio-Visual Branchformer
The model architecture previously described, composed of two independent modality-specific encoders, corresponds to the most prevalent approach adopted in the context of AVSR [10, 11, 45, 14, 13]. While recent works have explored unified encoder architectures [12, 19, 20], determining the optimal cross-modal architecture remains an ongoing challenge. In this paper, unlike prior works, we introduce a novel framework that, to the best of our knowledge, constitutes the first attempt to harness the flexibility and interpretability offered by encoder architectures, such as the Branchformer [21], in the design of parameter-efficient AVSR systems. The proposed method comprises multiple steps.
Training Modality-Specific Models. We first estimated models for audio-only and video-only inputs separately. Each model used a modality-specific Branchformer encoder. Implementation details can be found in Subsection IV-C.
Analysing Branchformer Encoder Layers. As described in Subsection III-A, the Branchformer encoder allowed us to analyse which branch was more relevant at each encoder layer for each of our estimated modality-specific models. Figure 2 reflects how the encoder adopted a different architecture depending on the addressed modality for the LRS2-BBC dataset. The increased reliance on global context relationships in the video-only scenario, compared to the audio-only setting, could be attributed to the absence of visible cues from lip movements when interpreting speech. Similar behaviours were observed across the datasets explored in our work.
@@ III-B The Tailored Audio-Visual Branchformer
Designing a Tailored AVSR Model. We also observed that, in most cases and regardless the modality, one of the relative branch weights of the modality-specific models tended to approach saturation, with values consistently close to 1.0. Our hypothesis was that, due to this saturation, the branch associated with the lower impact could be removed without a significant decrease in model performance. Based on this analysis, we designed a tailored, unified, parameter-efficient encoder for AVSR, whose overview scheme is depicted in Figure 3. This tailored design involves the selection of ad-hoc global- or local-context processing modules depending on the modality and the specific encoder layer level we are addressing. Specifically, our proposed tailored encoder is composed of two parallel branches but, in this case, each one is in charge of processing a different modality. Each layer of these branches is defined by a self-attention or cgMLP module depending on which of them was more relevant in its corresponding modality-specific model. Unlike the original Branchformer encoder, we do not merge the encoded representation of each branch. Doing so would fundamentally alter the nature of the audio- and video-only representations, thus hindering the ad-hoc design described above by not having a base model to select one module focused on either global or local context relationships. Nonetheless, we share the two feed-forward modules with the intention of enhancing the model’s cross-modal capabilities. Consequently, in order to facilitate the model to distinguish between both modalities, we introduced a modality embedding that condition the frontend speech representations. As supported by our extensive experiments, this tailored architecture design allowed us to reduce the number of model parameters by nearly half compared to conventional AVSR systems, while achieving state-of-the-art recognition rates.
@@ IV Experiments
The models proposed in this work were implemented using the open-source ESPNet toolkit [46]. Experiments were conducted on a 24-core 3.40MHz Intel i7-13700K CPU and a GeForce RTX 4090 GPU with 24GB memory.
IV-A Databases
@@
LRS2-BBC [10] is a large-scale English database composed of around 224 hours collected from BBC TV programs. It consists of a pre-training set with 96,318 samples (195 hours), a training set with 45,839 samples (28 hours), a validation set with 1,082 samples (0.6 hours), and a test set with 1,243 samples (0.5 hours). It offers more than 2M running words with a vocabulary size of around 40k.
LRS3-TED. [22] is the largest publicly audio-visual English database, providing around 438 hours collected from TED talks. It consists of a ‘pre-train’ set with 118,516 samples (407 hours), a ‘train-val’ set with 31,982 samples (30 hours), and a test set with 1,321 samples (0.9 hours). It offers more than 4M running words with a vocabulary size of around 50k.
MuAViCes [23] is a large-scale multilingual corpus collected from TED and TEDx talks. In our work, we only considered its Spanish partition, offering around 181 hours of data. It is split into a training set with 102,035 samples (177.5 hours), a validation set with 906 samples (1.6 hours), and a test set of 1012 samples (1.7 hours). It comprises around 1.7M running words with a vocabulary size of around 58k.
CMU-MOSEASes [47] is a large-scale database collected from YouTube monologue videos covering 4 different languages. In our work, we only considered its Spanish partition. According to [6], we defined a training set with 8,253 samples (15.7 hours) and a test set with 329 samples (0.6 hours). The database provides more than 150k running words with a vocabulary size of around 14k.
LIP-RTVE [48] is a challenging Spanish database collected from TV newscast programs, providing around 13 hours of data. Its speaker-independent partition consists of a training set with 7,142 samples (9 hours), a validation set with 1638 samples (2 hours), and a test set with 1572 samples (2 hours). It comprises around 300 speakers and more than 100k running words with a vocabulary size of around 10k.
VLRF [49] is a Spanish database primarily designed to study the feasibility of the VSR task, providing 1 hour of data. It was recorded in controlled laboratory settings, providing a speaker-dependent partition composed of 24 speakers, each one with 25 assigned unrelated but predefined sentences. It is split into a training set with 480 samples (0.92 hours) and a test set with 120 samples (0.23 hours), comprising around 4k running words with a vocabulary size of 1374.
@@
Audio- and Video-Only. These modality-specific systems constitute the baseline from which we will extract the Branchformer-based knowledge to design our proposed tailored AVSR system. Note that in these audio- and video-only settings, the model architecture described in Subsection III-A is modified so that, in addition to removing the audio-visual fusion module, only the corresponding fronted is considered. The audio- and video-only models comprise around 51.2M and 60.7M parameters, respectively, for both languages.
@@ IV-C Implementation Details
TABLE I: Comparison of WER (%) to the State of the Art on the Test Sets of the Datasets Composing the English Benchmark for Audio-Only (A), Video-Only (V), and Audio-Visual (AV) Settings. Works Using Non-Publicly Resources Were Not Considered.
Method Modality No. of Parameters Total Hours Additional Datasets Dataset
LRS2-BBC LRS3-TED
Ma et al. [11] A $\sim$100M 381 / 595 LRW 3.9 2.3
Shi et al. [12] A 325M 2,192 VoxCeleb2 - 1.5
Ma et al. [14] A 250.4M 818 LRW 2.6 1.5
Ma et al. [14] A 250.4M 3,448 LRW, VoxCeleb2, AVSpeech 1.5 1.0
Burchi and Timofte [13] A 61.7M 818 LRW 2.4 2.0
Burchi and Timofte [50] A 119M 3,116 LRW, VoxCeleb2, AVSpeech - 0.7
Ours (Conventional) A 51.2M 818 LRW 3.9$\pm$0.5 2.4$\pm$0.4
Ours (Tailored) A 43.3M 818 LRW 4.1$\pm$0.5 2.6$\pm$0.4
Ma et al. [11] V $\sim$100M 381 / 595 LRW 37.9 43.3
Prajwal et al. [45] V - 698 - 28.9 40.6
Shi et al. [12] V 325M 2,192 VoxCeleb2 - 26.9
Ma et al. [6] V 52M 818 LRW 27.3 34.5
Liu et al. [51] V >250M 6720 LRW, VoxCeleb2, AVSpeech, Synthetic - 16.9
Ma et al. [14] V 250.4M 818 LRW 27.9 33.0
Ma et al. [14] V 250.4M 3,448 LRW, VoxCeleb2, AVSpeech 14.6 19.1
Burchi and Timofte [13] V 61.7M 818 LRW 29.8 37.5
Burchi and Timofte [50] V 130M 3,116 LRW, VoxCeleb2, AVSpeech - 25.5
Ours (Conventional) V 60.7M 818 LRW 29.9$\pm$1.6 37.0$\pm$1.6
Ours (Tailored) V 51.3M 818 LRW 33.6$\pm$1.7 41.2$\pm$1.7
Ma et al. [11] AV $\sim$100M 381 / 595 LRW 3.7 2.3
Shi et al. [12] AV 325M 2,192 VoxCeleb2 - 1.3
Ma et al. [14] AV 250.4M 3,448 LRW, VoxCeleb2, AVSpeech 1.5 0.9
Burchi and Timofte [13] AV 61.7M 818 LRW 2.3 1.8
Burchi and Timofte [50] AV 197M 3,116 LRW, VoxCeleb2, AVSpeech - 0.8
Ours (Conventional) AV 103.5M 818 LRW 3.0$\pm$0.4 2.4$\pm$0.4
Ours (Tailored) AV 59.3M 818 LRW 3.0$\pm$0.4 2.1$\pm$0.4
TABLE II: Comparison of WER (%) to the State of the Art on the Test Sets of the Datasets Composing the Spanish Benchmark for Audio-Only (A), Video-Only (V), and Audio-Visual (AV) Settings.
Method Modality No. of Parameters Total Hours Additional Datasets %WER
MuAViCes
Anwar et al. [23] A 325M 2,370 LRW, LRS3, VoxCeleb2 16.5
Burchi et al. [50] A 119M 4,957 LRW, LRS3, VoxCeleb2, AVSpeech, MuAViC 7.9
[…]
Ours (Tailored) A 43.3M 203 - 12.8$\pm$0.8
Ma et al. [6] V 52M 1,546 LRW, LRS2&3, AVSpeech 56.6$\pm$0.3
Kim et al. [25] V 325M 2,825 LRW, LRS3, mTEDx, MLS 70.2
Yeo et al. [52] V 250.4M 3,832 LRW, LRS2&3, VoxCeleb2, AVSpeech 45.7
[…]
Ours (Tailored) V 51.3M 1,021 LRW, LRS2&3 80.9$\pm$1.6
Anwar et al. [23] AV 325M 2,370 LRW, LRS3, VoxCeleb2 15.9
Hong et al. [26] AV $>$325M 2,900 LRW, LRS3, VoxCeleb2, MuAViC 18.0
Burchi et al. [50] AV 197M 4,957 LRW, LRS3, VoxCeleb2, AVSpeech, MuAViC 8.2
[…]
Ours (Tailored) AV 58.3M 361 LRW 14.3$\pm$1.1
CMU-MOSEASes
Ma et al. [6] A 50M 1,546 LRW, LRS2&3, AVSpeech 15.4$\pm$0.1
[…]
Ours (Tailored) A 43.3M 203 - 15.1$\pm$1.3
Ma et al. [6] V 52M 1,546 LRW, LRS2&3, AVSpeech 44.6$\pm$0.6
[…]
Ours (Tailored) V 51.3M 1,021 LRW, LRS2&3 52.6$\pm$2.2
Ours (Conventional) AV 103.5M 361 LRW 13.0$\pm$1.3
Ours (Tailored) AV 58.3M 361 LRW 15.3$\pm$1.4
LIP-RTVE
Ours (Prior Work) [48] A - 9 - 15.3$\pm$0.8
[…]
Ours (Tailored) A 43.3M 203 - 8.9$\pm$0.5
Ours (Prior Work) [48] V - 9 - 95.9$\pm$0.2
[…]
Ours (Tailored) V 51.3M 1,021 LRW, LRS2&3 64.0$\pm$1.1
Ours (Conventional) AV 103.5M 361 LRW 6.7$\pm$0.4
Ours (Tailored) AV 58.3M 361 LRW 7.6$\pm$0.5
VLRF
Ours (Prior Work) [53] A - 1 - 9.1$\pm$3.0
[…]
Ours (Tailored) A 43.3M 203 - 6.0$\pm$2.3
Fernández-López [54] V - - Synthetic 72.9
Ours (Prior Work) [53] V - 1 - 59.7$\pm$4.3
[…]
Ours (Tailored) V 51.3M 1,021 LRW, LRS2&3 23.0$\pm$3.2
Ours (Conventional) AV 103.5M 361 LRW 5.0$\pm$1.7
Ours (Tailored) AV 58.3M 361 LRW 5.8$\pm$2.2
  • $\dagger$
    Though results were reported for mTEDx [55], it corresponds to MuAViC, as noted by Anwar et al. [23].
@@
Video Processing. All videos were first sampled at 25 fps. Similar to [11, 6, 13], we cropped patches of 96$\times$96 pixels centered on the speaker’s mouth. We used the RetinaFace face detector [56] and the Face Alignment Network [57] to detect 68 facial landmarks. Subsequently, we applied a similarity transformation w.r.t. a neutral reference frame, thereby removing possible translation and scaling variations. The cropped lip regions were converted into gray-scale, and normalized based on the overall mean and variance of the training set.
@@
Text Processing. We first normalize texts by punctuation removal and uppercasing. Similar to [6], we then used a character-level tokenizer with a vocab size of 41 and 37 for English and Spanish, respectively. In both cases, special symbols, such as the ‘space’, ’end of sentence’, and the ‘blank’ tokens, were included.
@@
Data Augmentation. We applied SpecAugment [58] on our audio cues during training with two frequency masks with mask size parameter F = 27 and five time masks with adaptive size pS = 0.05. Furthermore, we introduce speed perturbation by randomly selecting a scale factor between 0.9 and 1.1 from a uniform distribution. Regarding video, random cropping of 88$\times$88 pixels, horizontal flipping, and time masking [6] with a maximum of 0.4 seconds were applied during training.
@@
Training Settings. Similar to [13], we used the Adam optimizer [59] and a Noam scheduler [24] with 10k warmup steps, yielding a peak learning rate of 0.001 with a batch size of 64 samples. All the proposed models were trained for 100 epochs. Those utterances that were longer than 20 seconds were discarded. In all cases, the loss weight $\alpha$ in Eq. (5) was set to 0.1. Similar to most works in the field, the visual frontend was initialized in all our experiments with the publicly released weights [^1] https://github.com/mpc001/Lipreading_using_Temporal_Convolutional_Networks trained on the word-level LRW classification task [60]. Additionally, encountering convergence issues with the Spanish video-only system and inspired by the work carried out by Ma et al. [6], we adopted a cross-lingual learning strategy, initializing both the visual frontend and the encoder backbone with the weights estimated for the English benchmark. Subsequently, we conducted fine-tuning following the training settings outlined in [6].
@@
Inference Settings. Similar to [6], we set the LM weight $\beta$ of Eq. (6) to 0.6 and 0.4 for English and Spanish, respectively. For English, we set a word insertion penalty of 0.5 and a beam size of 40, while for Spanish, we set the word insertion penalty and the beam size to 0.0 and 30, respectively. Regarding the CTC/Attention balance, we set the weight $\lambda$ of Eq. (6) to 0.1. Note that, during inference, we averaged the model parameters on the 10 epochs with the top validation performance.
@@
Evaluation Metric. Reported results were evaluated by the well-known Word Error Rate (WER) with 95% confidence intervals using the bootstrap method described by [61].
@@
Limitations. Numerous studies [10, 11, 12, 13] investigated the benefits of techniques involving either masking or incorporating noise into acoustic cues during training. While these methods have demonstrated improvements in the robustness of AVSR systems, we consciously refrained from adopting these strategies. Our rationale was to properly study the cross-modality learning capabilities and noise robustness of our proposed method strictly from an architectural design perspective, without incorporating such data augmentation artifacts.

### v3

@@ III-B The Tailored Audio-Visual Branchformer
Designing a Tailored AVSR Model. We also observed that, in most cases and regardless of the modality, one of the relative branch weights of the modality-specific models tended to approach saturation, with values consistently close to 1. Our hypothesis was that, due to this saturation, the branch associated with the lower impact could be removed without a significant decrease in model performance. Based on this analysis, we designed a tailored, unified, parameter-efficient encoder for AVSR, whose overview scheme is depicted in Figure 3. This tailored design involves the selection of ad-hoc global- or local-context processing modules depending on the modality and the specific encoder layer level we are addressing. Specifically, our proposed tailored encoder is composed of 12 layers with two parallel branches but, in this case, each one is in charge of processing a different modality. Each layer of these branches is defined by a self-attention or cgMLP module depending on which of them was more relevant in its corresponding modality-specific model. Formally, this can be expressed as follows:
$\delta_{im}(x_{m})=\begin{cases}cgMLP(x_{m})&\text{if }E_{im}^{global}(x_{m})<E_{im}^{local}(x_{m}),\\ SelfAttention(x_{m})&\text{otherwise,}\end{cases}$ (7)
where $i$ denotes the layer index, $m\in\{audio,video\}$ indicates the modality, and $x_{m}$ represents the input features for modality $m$. Therefore, $\delta_{im}$ corresponds to the context-specific processing module of the $i$-th layer of the tailored, unified audio-visual encoder. Additionally, $E_{im}^{global}(x_{m})$ and $E_{im}^{local}(x_{m})$ refer to the scores provided by our modality-specific models that indicate the relevance of global and local context at layer $i$ for modality $m$, respectively.
Unlike the original Branchformer encoder, we do not merge the encoded representation of each branch. Doing so would fundamentally alter the nature of the audio- and video-only representations, thus hindering the ad-hoc design described above by not having a base model to select one module focused on either global or local context relationships. Nonetheless, we share the two feed-forward modules with the intention of enhancing the model’s cross-modal capabilities. Consequently, in order to facilitate the model to distinguish between both modalities, we introduced a modality embedding that conditions the frontend speech representations. A preliminary ablation study showed that, although no significant differences were observed, incorporating the modality embedding consistently led to slight improvements.
As supported by our extensive experiments, this tailored architecture design allowed us to reduce the number of model parameters by nearly half compared to conventional AVSR systems, while achieving state-of-the-art recognition rates across diverse benchmarks.
@@ IV Experiments
The models proposed in this work were implemented using the open-source ESPNet toolkit [watanabe18espnet]. Experiments were conducted on a 24-core 3.40MHz Intel i7-13700K CPU and a GeForce RTX 4090 GPU with 24GB memory.
IV-A Datasets
@@
LRS2-BBC [afouras2018deep] is a large-scale English database composed of around 224 hours collected from BBC TV programs. It consists of a pre-training set with 96,318 samples (195 hours), a training set with 45,839 samples (28 hours), a validation set with 1,082 samples (0.6 hours), and a test set with 1,243 samples (0.5 hours). It offers more than 2M running words with a vocabulary size of around 40k.
LRS3-TED [afouras2018lrs3] is the largest publicly audio-visual English database, providing around 438 hours collected from TED talks. It consists of a ‘pre-train’ set with 118,516 samples (407 hours), a ‘train-val’ set with 31,982 samples (30 hours), and a test set with 1,321 samples (0.9 hours). It offers more than 4M running words with a vocabulary size of around 50k.
MuAViCes [anwar23muavic] is a large-scale multilingual corpus collected from TED and TEDx talks. In our work, we only considered its Spanish partition, offering around 181 hours of data. It is split into a training set with 102,035 samples (177.5 hours), a validation set with 906 samples (1.6 hours), and a test set of 1012 samples (1.7 hours). It comprises around 1.7M running words with a vocabulary size of around 58k.
CMU-MOSEASes [zadeh2020moseas] is a large-scale database collected from YouTube monologue videos covering 4 different languages. In our work, we only considered its Spanish partition. According to [ma2022visual], we defined a training set with 8,253 samples (15.7 hours) and a test set with 329 samples (0.6 hours). The database provides more than 150k running words with a vocabulary size of around 14k.
LIP-RTVE [lrec2022liprtve] is a challenging Spanish database collected from TV newscast programs, providing around 13 hours of data. Its speaker-independent partition consists of a training set with 7,142 samples (9 hours), a validation set with 1638 samples (2 hours), and a test set with 1572 samples (2 hours). It comprises around 300 speakers and more than 100k running words with a vocabulary size of around 10k.
VLRF [fernandez2017towards] is a Spanish database primarily designed to study the feasibility of the VSR task, providing 1 hour of data. It was recorded in controlled laboratory settings, providing a speaker-dependent partition composed of 24 speakers, each one with 25 assigned unrelated but predefined sentences. It is split into a training set with 480 samples (0.92 hours) and a test set with 120 samples (0.23 hours), comprising around 4k running words with a vocabulary size of 1374.
@@
Audio- and Video-Only. These modality-specific systems constitute the baseline from which we will extract the Branchformer-based knowledge to design our proposed tailored AVSR system. Note that in these audio- and video-only settings, the model architecture described in Subsection III-A is modified so that, in addition to removing the audio-visual fusion module, only the corresponding frontend is considered. The audio- and video-only models comprise around 51.2M and 60.7M parameters, respectively, for both languages.
@@ IV-C Implementation Details
Video Processing. All videos were first sampled at 25 fps. Similar to [ma2021conformers, ma2022visual, burchi2023audio], we cropped patches of 96$\times$96 pixels centered on the speaker’s mouth. We used the RetinaFace face detector [deng2020retina] and the Face Alignment Network [bulat2017facealign] to detect 68 facial landmarks. Subsequently, we applied a similarity transformation w.r.t. a neutral reference frame, thereby removing possible translation and scaling variations. The cropped lip regions were converted into gray-scale, and normalized based on the overall mean and variance of the training set.
@@
Text Processing. We first normalize texts by punctuation removal and uppercasing. Similar to [ma2022visual], we then used a character-level tokenizer with a vocab size of 41 and 37 for English and Spanish, respectively. In both cases, special symbols, such as the ‘space’, ’end of sentence’, and the ‘blank’ tokens, were included.
@@
Data Augmentation. We applied SpecAugment [park2020specaugment] on our audio cues during training with two frequency masks with mask size parameter F = 27 and five time masks with adaptive size pS = 0.05. Furthermore, we introduce speed perturbation by randomly selecting a scale factor between 0.9 and 1.1 from a uniform distribution. Regarding video, random cropping of 88$\times$88 pixels, horizontal flipping, and time masking [ma2022visual] with a maximum of 0.4 seconds were applied during training.
@@
Training Settings. Similar to [burchi2023audio], we used the Adam optimizer [kingma2014adam] and a Noam scheduler [vaswani2017attention] with 10k warmup steps, yielding a peak learning rate of 0.001 with a batch size of 64 samples. All the proposed models were trained for 100 epochs. Those utterances that were longer than 20 seconds were discarded. In all cases, the loss weight $\alpha$ in Eq. (5) was set to 0.1. Similar to most works in the field, the visual frontend was initialized in all our experiments with the publicly released weights [^1] https://github.com/mpc001/Lipreading_using_Temporal_Convolutional_Networks trained on the word-level LRW classification task [ma2021towards]. Additionally, encountering convergence issues with the Spanish video-only system and inspired by the work carried out by Ma et al. [ma2022visual], we adopted a cross-lingual learning strategy, initializing both the visual frontend and the encoder backbone with the weights estimated for the English benchmark. Subsequently, we conducted fine-tuning following the training settings outlined in [ma2022visual].
@@
Inference Settings. Similar to [ma2022visual], we set the LM weight $\beta$ of Eq. (6) to 0.6 and 0.4 for English and Spanish, respectively. For English, we set a word insertion penalty of 0.5 and a beam size of 40, while for Spanish, we set the word insertion penalty and the beam size to 0.0 and 30, respectively. Regarding the CTC/Attention balance, we set the weight $\lambda$ of Eq. (6) to 0.1. Note that, during inference, we averaged the model parameters on the 10 epochs with the top validation performance.
@@
Evaluation Metric. Reported results were evaluated by the well-known Word Error Rate (WER) with 95% confidence intervals using the bootstrap method described by [bisani2004bootstrap].
@@
Limitations. Numerous studies [afouras2018deep, ma2021conformers, shi2022learning, burchi2023audio] investigated the benefits of techniques involving either masking or incorporating noise into acoustic cues during training. While these methods have demonstrated improvements in the robustness of AVSR systems, we consciously refrained from adopting these strategies. Our rationale was to properly study the cross-modality learning capabilities and noise robustness of our proposed method strictly from an architectural design perspective, without incorporating such data augmentation artifacts.